Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question

Expert replies
by yuvarajait » Sat Apr 16, 2011 10:56 pm
A medical researcher must choose one of 14 patients to receive an experimental
medicine called Progaine. The researcher must then choose one of
the remaining 13 patients to receive another medicine, called Ropecia.
Finally, the researcher administers a placebo to one of the remaining 12
patients. All choices are equally random. If Donald is one of the 14 patients,
what is the probability that Donald receives either Progaine or Ropecia?

The answer given is 1/14+1/14 = 1/7, which i feel as incorrect, since the chances of getting ropecia is only 1/13. Could anyone help understanding this. Thanks!
Join the discussion
Source: — Problem Solving |

by manpsingh87 » Sun Apr 17, 2011 12:37 am
yuvarajait wrote:A medical researcher must choose one of 14 patients to receive an experimental
medicine called Progaine. The researcher must then choose one of
the remaining 13 patients to receive another medicine, called Ropecia.
Finally, the researcher administers a placebo to one of the remaining 12
patients. All choices are equally random. If Donald is one of the 14 patients,
what is the probability that Donald receives either Progaine or Ropecia?

The answer given is 1/14+1/14 = 1/7, which i feel as incorrect, since the chances of getting ropecia is only 1/13. Could anyone help understanding this. Thanks!
well my advice is don't get confused with the wording,each patient has an equal chance of being selected for any of the three experimental medicine..!!!
O Excellence... my search for you is on... you can be far.. but not beyond my reach!
Join the discussion

by HSPA » Sun Apr 17, 2011 12:48 am
A) selected first = 1/14
B) getting selected for the second = 13/14* 1/13 = 1/14
Here 13/14 stands for not getting selected for first medicine Progaine and the next 1/13 is getting selected for Ropecia.

P(AUB) = P(A) + P(B) = 1/14+ 1/14 = 1/7
First take: 640 (50M, 27V) - RC needs 300% improvement
Second take: coming soon..
Regards,
HSPA.
Join the discussion

by Brent@GMATPrepNow » Sun Apr 17, 2011 6:59 am
yuvarajait wrote:A medical researcher must choose one of 14 patients to receive an experimental
medicine called Progaine. The researcher must then choose one of
the remaining 13 patients to receive another medicine, called Ropecia.
Finally, the researcher administers a placebo to one of the remaining 12
patients. All choices are equally random. If Donald is one of the 14 patients,
what is the probability that Donald receives either Progaine or Ropecia?

The answer given is 1/14+1/14 = 1/7, which i feel as incorrect, since the chances of getting ropecia is only 1/13. Could anyone help understanding this. Thanks!
You can also solve the question using the complement.
P(Progaine or Ropecia) = 1 - P(neither Progaine nor Ropecia)

P(neither Progaine nor Ropecia) = P(not getting selected for Progaine AND not getting selected for Ropecia)
= P(not getting selected for Progaine) x P(not getting selected for Ropecia)
= (13/14) x (12/13)
= 12/14
= 6/7

So, P(Progaine or Ropecia) = 1 - P(neither Progaine nor Ropecia)
= 1 - 6/7
= 1/7
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Ian Stewart » Sun Apr 17, 2011 11:13 am
yuvarajait wrote:A medical researcher must choose one of 14 patients to receive an experimental
medicine called Progaine. The researcher must then choose one of
the remaining 13 patients to receive another medicine, called Ropecia.
Finally, the researcher administers a placebo to one of the remaining 12
patients. All choices are equally random. If Donald is one of the 14 patients,
what is the probability that Donald receives either Progaine or Ropecia?

The answer given is 1/14+1/14 = 1/7, which i feel as incorrect, since the chances of getting ropecia is only 1/13. Could anyone help understanding this. Thanks!
I posted about a similar problem a moment ago, but you can also just imagine lining up your 14 people, giving the 1st person in line Progaine, the 2nd person Ropecia, and the 3rd person the placebo. The question then becomes 'if you line up 14 people, including Don, what is the probability Don is one of the first 2 people in line?' There are 14 places to put Don, 2 of which are in the first two positions in line, so the answer is 2/14 = 1/7.

Note that the chance of getting Ropecia is *not* 1/13; it is still 1/14. It is only 1/13 if you *know* you were not selected to receive Progaine.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion